OpenStudy (anonymous):

Solve rational equations (k-2)/(k+1) =(7k+22)/(k^2+6k+5)

OpenStudy (ranga):

cross multiply and simplify.

OpenStudy (anonymous):

Don't you use the quadratic formula?

OpenStudy (ranga):

Well, you can solve this a couple of ways. You can factor the (k^2+6k+5) first and see if any factor cancels out. Let us try that approach. What are the factors of (k^2+6k+5)?

OpenStudy (anonymous):

k+5 and k+1

OpenStudy (ranga):

Exactly! So the (k+1) in the denominator on the right cancels out with the (k + 1) in the denominator of the left. Now you can simplify.

OpenStudy (anonymous):

Im stuck with the 7k+22

OpenStudy (anonymous):

Im suppose to get the answer k=-4 and k=8

OpenStudy (ranga):

(k-2)/(k+1) = (7k+22)/(k^2+6k+5) (k-2)/(k+1) = (7k+22)/((k+5)(k+1)) (k+1) cancels out. (k-2) = (7k+22)/(k+5) cross multiply (k-2) * (k + 5) = 7k + 22 k^2 + 5k - 2k - 10 = 7k + 22 k^2 + 3k - 10 - 7k - 22 = 0 k^2 - 4k - 32 = 0 factor and solve for k.

OpenStudy (anonymous):

Thanks for typing all of that! Then you can use the quadratic equation?

OpenStudy (ranga):

You can use the quadratic equation. But my preference is to always try to see if I can factor it. k^2 - 4k - 32 = 0 write it as: k^2 - 8k + 4k - 32 = 0 k(k - 8) + 4(k - 8) = 0 (k + 4)(k - 8) = 0 k = -4 or 8

OpenStudy (anonymous):

I see! It would be faster with factoring now! Thank you so much for walking me through this!

OpenStudy (ranga):

You are welcome. BTW, I answered your other question with long division although it is closed now.

OpenStudy (anonymous):

Oh! HA! I figured it out! Thank you

OpenStudy (anonymous):

I do have another one that Im going to post.

OpenStudy (anonymous):

f(x)=\[\sqrt[7]{x+20?}\] a. find f(-60) b. f(x) =-3

OpenStudy (ranga):

For f(-60) you just need to put x = -60 into the function and evaluate.

OpenStudy (anonymous):

I did that but got stuck with the 7. Im going over my notes now to try and figure it out.

OpenStudy (ranga):

the little 7 outside the radical means the seventh root. It is the same as raising it to ^(1/7)

OpenStudy (ranga):

\[\Large f(x) = \sqrt[7]{(x + 20)} = (x + 20)^{\frac{ 1 }{ 7}}\]

OpenStudy (anonymous):

Im suppose to have the answer f(-60) aproxi -1.694

OpenStudy (ranga):

Yes. f(x) = (x + 20)^(1/7) Find f(-60). Put x = -60 f(-60) = (-60 + 20)^(1/7) = (-40)^(1/7) Use your calculator to find (-40)^(1/7) If you don't have a scientific calculator, you can use Google calculator.

OpenStudy (anonymous):

The index messes me up.

OpenStudy (ranga):

That is why rewriting the seventh root as exponent (1/7) will make it easier to understand.

OpenStudy (anonymous):

I do have a scientific calc and I will check it out. Thank you!

OpenStudy (anonymous):

Yes, it does help! Can you do that will all of them?

OpenStudy (ranga):

Yes. BTW, I tried Google calc and I am getting -1.6938

OpenStudy (anonymous):

OhhH! I got it1 Yeah!

OpenStudy (ranga):

Alright! Now on to part b.

OpenStudy (anonymous):

Are you a professor or you just love math?

OpenStudy (ranga):

Not a prof. Just volunteering cuz it is a good workout for the gray matter and at the same time it gives the satisfaction of helping.

OpenStudy (anonymous):

I know how that is... I decided to go back to school after a long time out and Math is my worst subject! But I'm making it. Thank so much for being on here. B. f(x)=-3

OpenStudy (ranga):

No problem. f(x) = (x + 20)^(1/7) f(x) = -3 So we need to solve: -3 = (x + 20)^(1/7) We need to get rid of the exponent 1/7. To do that raise both sides to the power of 7. 7 and 1/7 will cancel out on the right.

OpenStudy (ranga):

(-3)^7 = { (x + 20)^(1/7) }^7 = x + 20 (-3)^7 = x + 20 to isolate x, add -20 to both sides: (-3)^7 - 20 = x you can use your calc to find (-3)^7 - 20

OpenStudy (ranga):

I am getting x = -2207.

OpenStudy (anonymous):

Yes!!! Great thank you! I am now working it out to understand better.

OpenStudy (anonymous):

You have have been a huge help!

OpenStudy (ranga):

Glad to be able to assist. I am leaving this page now. Since it is a closed question the notification may not work to let me know if you had posted any reply or had any question. Even when the question is open the notification is not always timely. Do well in your math tests. Good luck.

OpenStudy (anonymous):

Thank you, I was finished with the questions. Thanks!